Approximately determine the number of cells in the human body. Model the human body with a cylinder 1.9 m high and with a radius of 0.1 m at its base. Model the cell with a cube with an edge length of 10µm.


1
Expert's answer
2021-03-04T11:57:20-0500

Let's first find the volume of the cylinder:


"V_{c}=\pi r^2h=\pi\cdot(0.1\ m)^2\cdot1.9\ m=0.06\ m^3."

Let's find the volume of the cell:


"V_{cell}=a^3=(10\cdot10^{-6}\ m)^3=1.0\cdot10^{-15}\ m^3."

Finally, we can determine the number of cells in the human body:


"N=\dfrac{V_c}{V_{cell}}=\dfrac{0.06\ m^3}{1.0\cdot10^{-15}\ m^3}=6.0\cdot10^{13}\ cells."